# EE-LLM: Large-Scale Training and Inference of Early-Exit Large Language Models with 3D Parallelism

### v2

@@ B.1 Methodology
The maximum number of microbatches that can be inserted into Part 1 or 2 of the explicit bubbles, without increasing training time per iteration, is $\lfloor(p-1)b/(f+b)\rfloor=\lfloor(p-1)/(f/b+1)\rfloor$, where $f/b$ is (an estimate of) the ratio between forward and backward time. To see this for Part 1 (resp. 2), one simply need to notice that the first (resp. last) pipeline stage has an bubble size $b(p-1)$, while the total forward and backward time for each microbatch is $f+b$. Dividing the first value by the second one concludes the proof.
@@
With $K$ inserted microbatches, the $i$-th microbatch is supposed to go through the forward pass of the first $K+1-i$ stages. However, if there is no early exit on Stage $K+1-i$, then this microbatch only need to go through the forward pass up to the last stage (among the first $K+1-i$ stages) that has at least one early exit.
@@
(1) There are some limitations to the proposed approach. For example, it requires that there is no tied/shared model parameter across pipeline stages; otherwise the gradients from partial forward/backward passes might (in theory) be harmful rather than beneficial for training the model. Another concern is that inserting microbatches into Part 1 of the explicit bubbles can cause additional overhead of activation memory in early stages. (2) While Part 1 of this approach requires the existence of early exits, Part 2 is actually applicable to training standard models without early exits as well. (3) One might raise concerns about the inefficiency of data usage, as some microbatches only undergo partial forward/backward passes. In general, this is not an issue for LLM pre-training, since training data is usually abundant, and the complete training process might not even go through one epoch of the data.
@@ B.2 Analysis
Claim 1 (Informal).
Under the above assumptions, the accumulated gradient of one training iteration, with extra updates from additional microbatches inserted into Part 2, remains (after some appropriate entrywise scaling) an unbiased estimate, but with reduced variance, of the gradient for the targeted population risk. A similar claim holds true for Part 1 as well, except for certain extreme (and unlikely) situations where gradients of different early/final losses have strong negative correlation.
@@
We start with analysis for Part 2, which is slightly simpler. Suppose that an additional microbatch goes through the full forward pass, followed by a partial backward pass covering only the last two stages, namely $\bm{\theta}_{4}$ and $\bm{\theta}_{3}$. Then, the gradient from this microbatch is
@@
here, the first equality is due to the partial backward pass, and the second equality follows from the observation that $\partial{\mathcal{L}}_{i}/\partial\bm{\theta}_{j}=\bm{0}$ for any $i<j$.
@@
We claim that this leads to reduced gradient variance, unless gradients of early and later losses have strong negative correlation. In the formal analysis of variance below, we consider scalars for notational simplicity, though it can be easily generalized to vectors or tensors by replacing scalar multiplication with inner product.
Proposition 2.
@@
[…]
@@
[…]

### v3

@@ C.2.1 Methodology
The maximum number of microbatches that can be inserted into Part 1 or 2 of the explicit bubbles, without increasing training time per iteration, is $\lfloor(p-1)b/(f+b)\rfloor=\lfloor(p-1)/(f/b+1)\rfloor$, where $f/b$ is (an estimate of) the ratio between forward and backward time. To see this for Part 1 (resp. 2), one simply needs to notice that the first (resp. last) pipeline stage has a bubble size $b(p-1)$, while the total forward and backward time for each microbatch is $f+b$. Dividing the first value by the second one concludes the proof.
@@
With $K$ inserted microbatches, the $i$-th microbatch is supposed to go through the forward pass of the first $K+1-i$ stages. However, if there is no early exit on Stage $K+1-i$, then this microbatch only needs to go through the forward pass up to the last stage (among the first $K+1-i$ stages) that has at least one early exit.
@@
(1) There are some limitations to the proposed approach. For example, it requires that there is no tied/shared model parameter across pipeline stages; otherwise, the gradients from partial forward/backward passes might (in theory) be harmful rather than beneficial for training the model. Another concern is that inserting microbatches into Part 1 of the explicit bubbles can cause additional overhead of activation memory in early stages. (2) While Part 1 of this approach requires the existence of early exits, Part 2 is actually applicable to training standard models without early exits as well. (3) One might raise concerns about the inefficiency of data usage, as some microbatches only undergo partial forward/backward passes. In general, this is not an issue for LLM pre-training, since training data is usually abundant, and the complete training process might not even go through one epoch of the data.
@@ C.2.2 Analysis
Claim C.1 (Informal).
Under the above assumptions, the accumulated gradient of one training iteration, with extra updates from additional microbatches inserted into Part 2, remains (after some appropriate entrywise scaling) an unbiased estimate, but with reduced variance, of the gradient for the targeted population risk. A similar claim holds true for Part 1 as well, except for certain extreme (and unlikely) situations where gradients of different early/final losses have a strong negative correlation.
@@
We start with an analysis for Part 2, which is slightly simpler. Suppose that an additional microbatch goes through the full forward pass, followed by a partial backward pass covering only the last two stages, namely $\bm{\theta}_{4}$ and $\bm{\theta}_{3}$. Then, the gradient from this microbatch is
@@
here, the first equality is due to Proposition 3.1 and the partial backward pass, while the second equality follows from the observation that $\partial{\mathcal{L}}_{i}/\partial\bm{\theta}_{j}=\bm{0}$ for any $i<j$.
@@
We claim that this leads to reduced gradient variance, unless gradients of early and later losses have a strong negative correlation. In the formal analysis of variance below, we consider scalars for notational simplicity, though it can be easily generalized to vectors or tensors by replacing scalar multiplication with inner product.
Proposition C.2.
@@
$\widehat{e}\coloneqq\widehat{a}_{N}+\widehat{b}_{N},\quad\widehat{e}_{+}\coloneqq\widehat{a}_{N+1}+\widehat{b}_{N},\quad\text{where}$
@@
${\mathbb{E}}[\widehat{e}]={\mathbb{E}}[\widehat{e}_{+}]=a^{\star}+b^{\star},\quad\text{var}(\widehat{e})-\text{var}(\widehat{e}_{+})=\frac{1}{N(N+1)}\text{var}(a_{1})+\frac{2}{N(N+1)}\text{cov}(a_{1},b_{1}).$
@@
Appendix D Implementations
The implementation of EE-LLM is based on Megatron-LM (Narayanan et al., 2021b), primarily extending Megatron-LM’s model architectures, pipeline scheduling, and inference service to support the training and inference of early-exit LLMs. We introduce each of these aspects in more detail below.
D.1 Model architectures
D.2 Pipeline scheduling
We have adjusted the existing 1F1B schedule for early-exit LLMs, as shown in Figure 3. To fill implicit bubbles and reduce GPU memory overhead, lazy loss functions of early-exit layers are returned together with outputs of the backbone network during forward steps. These lazy functions are not actually called until their corresponding auxiliary losses (cf. Section 3.1.1) are calculated in the backward steps. For the method of filling explicit bubbles proposed in Section 3.3, we have inserted partial forward/backward computation of additional microbatches into warm-up and cool-down phases of the 1F1B schedule. The number of inserted microbatches and partial forward/backward stages can be automatically calculated through the user-specified (estimate of) ratio between backward and forward time.
D.3 Inference service
For inference with pipeline parallelism, i.e. the pipeline-based approach proposed in Section 4, we have re-implemented the forward process. With our implementation, the first pipeline stage will wait for an exit signal from the early/final exits of all subsequent stages after its forward computation is completed. Each subsequent stage will send an exit signal and the output token to the first stage, if there is an exit within the stage that satisfies the exit condition. Upon receiving the signal and generated token, the first stage will immediately start the forward pass for generating the next token. With this implementation, regardless of the early-exit layers’ positions in subsequent stages, the inference service can immediately generate a token whenever early exiting happens on some stage, without waiting for the completion of the entire stage (except for the first stage).
For inference without pipeline parallelism, we have implemented a mechanism of KV recomputation, which is a variant of synchronized parallel decoding proposed recently in (Bae et al., 2023). In this approach, we maintain a list of the most recent tokens that have missing KV caches in deep layers due to early exiting. During each forward pass, we include these early-exit tokens in the current forward pass, which allows for direct recomputation of the KV caches for these tokens and thus avoids the issue of missing KV caches. Acceleration of sequence generation is still achieved, thanks to the batching effects of GPU computation. To avoid the endless accumulation of early-exit tokens, we enforce a full-model forward pass whenever the number of early-exit tokens reaches a pre-specified value.
Appendix E Preliminaries
Transformers.
The Transformer architecture (Vaswani et al., 2017; Tay et al., 2022) has been playing a dominant role in natural language processing (NLP) and large language models (LLMs) (Bommasani et al., 2021; Zhao et al., 2023a). It is typically composed of an input embedding layer, a stack of Transformer layers, and finally an output layer. Each Transformer layer consists of cross-attention and/or self-attention modules (Bahdanau et al., 2015; Kim et al., 2017; Parikh et al., 2016), a multi-layer perceptron (MLP), and layer normalization (LayerNorm (Ba et al., 2016) or RMSNorm (Zhang & Sennrich, 2019)). Transformers can be categorized into three types: encoder-only, encoder-decoder, and decoder-only. For the latter two, there is an output embedding matrix in the output layer, which transforms hidden states into logits on a (typically large) vocabulary that can be used for generating tokens. An LLM can be learned by unsupervised pre-training, e.g. minimizing the negative log-likelihood of next-token prediction on a large corpus (Radford et al., 2018, 2019). In this work, we focus on the decoder-only generative pre-training (GPT) Transformer architecture (Radford et al., 2018, 2019), though many of our ideas are widely applicable to other Transformer architectures or generic deep neural networks.
Early-exit LLMs.
An early-exit neural network can be obtained by adding to a standard neural network some early-exit layers that turn intermediate hidden states into early outputs (Xin et al., 2020; Schwartz et al., 2020). During inference for a given input, the model starts a forward pass and decides (at each early exit) whether to return an output or continue forwarding via certain rules, e.g. to return an output whenever the confidence of prediction is above a pre-defined threshold (Schwartz et al., 2020; Schuster et al., 2022).
The standard way of training an early-exit model is to minimize a weighted sum of early-exit and final-exit training losses (Schwartz et al., 2020; Schuster et al., 2022). Note that early-exit layers bring additional computational overhead to training. This is especially the case for LLMs, primarily due to the large output embedding matrix of size $h\times V$ within each early-exit layer, where $h$ is the hidden dimension and $V$ is the vocabulary size. We call an early-exit layer minimalistic if it has the same structure as the final output layer of the GPT model architecture (Radford et al., 2018, 2019), which includes an output embedding matrix, plus an optional LayerNorm/RMSNorm in front of it. Additional modules can be added to early-exit layers for increased expressivity and adaptivity of early exits.
3D parallelism.
3D parallelism refers to the combination of data, tensor, sequence and pipeline parallelism, and has been implemented in state-of-the-art LLM frameworks such as Megatron-LM (Shoeybi et al., 2019; Narayanan et al., 2021b; Korthikanti et al., 2022; Smith et al., 2022), DeepSpeed (Rasley et al., 2020; Smith et al., 2022), Mesh-TensorFlow (Shazeer et al., 2018), Alpa (Zheng et al., 2022), InternLM (Team, 2023), among others. With data parallelism, each GPU handles the forward and backward computation for one part of the data batch, and then the results are aggregated at the end of the training iteration. When the model is too large to fit in a single GPU, model partitioning becomes necessary and can be used in conjunction with data parallelism. With tensor (and sequence) parallelism, each large module (e.g. a linear layer) is divided into multiple pieces that are assigned to different GPUs, so that each computational task related to it (e.g. large matrix multiplication) can be divided into smaller tasks and solved in parallel. One major limitation of tensor (and sequence) parallelism is that it requires expensive collective communication such as all-reduce operations, and thus is only viable for high-end GPUs within the same computing node, with high-bandwidth communication among them.
Pipeline parallelism (Narayanan et al., 2019, 2021a; Fan et al., 2021; Li & Hoefler, 2021), on the other hand, partitions a deep model along the depth dimension into multiple pipeline stages. Moreover, each data batch is divided into multiple microbatches, and their forward/backward computational tasks are scheduled among those multiple pipeline stages. More specifically, each stage performs the forward computation for each microbatch and sends the resulting hidden states to another stage; later on, it performs the backward computation for the same microbatch after receiving the gradients of the training objective with respect to the sent hidden states. Pipeline parallelism only requires sparse and inexpensive point-to-point (P2P) communication between pipeline stages, which makes it applicable and oftentimes must-have in much broader scenarios when tensor (and sequence) parallelism is infeasible or insufficient, whether in GPU clusters or in decentralized settings (Yuan et al., 2022; Yang et al., 2023). The main concern with pipeline parallelism is its low utilization rate of computational resources, due to pipeline bubbles and load imbalance across pipeline stages (Narayanan et al., 2021b).
Appendix F Related works
Early exiting.
As introduced in Section 1, early exiting has been widely applied for accelerating inference of deep neural networks in the literature (Graves, 2016; Liu et al., 2020; Hou et al., 2020; Zhou et al., 2020; Elbayad et al., 2020; Schwartz et al., 2020; Xin et al., 2020; Schuster et al., 2021; Xin et al., 2021; Li et al., 2021; Hu et al., 2023; Teerapittayanon et al., 2016; Huang et al., 2018; Kaya et al., 2019; Laskaridis et al., 2021; Scardapane et al., 2020; Han et al., 2022; Xu & McAuley, 2023; Dai et al., 2023). In terms of early-exit Transformers in the NLP domain, the majority of prior works are focused on BERT (Devlin et al., 2019) or other encoder-only models for classification tasks (Liu et al., 2020; Hou et al., 2020; Zhou et al., 2020; Elbayad et al., 2020; Schwartz et al., 2020; Xin et al., 2020; Schuster et al., 2021; Xin et al., 2021; Li et al., 2021; Hu et al., 2023). Recent works have begun to study token-wise early exiting for accelerating inference of encoder-decoder or decoder-only LLMs in autoregressive sequence generation (Schuster et al., 2021; Corro et al., 2023; Bae et al., 2023; Varshney et al., 2023). While they are largely focused on designing inference mechanisms, the lack of support in prior works for training early-exit models with massive 3D parallelism inevitably poses an obstacle to truly scaling up early-exit LLMs. Our work on EE-LLM is one important step towards eliminating this obstacle, and we also contribute a novel pipeline-based inference mechanism along the way.
Early-exit inference with KV caching.
Several approaches have been recently proposed to resolve the conflict between early exiting and KV caching in autoregressive generation. One approach (Elbayad et al., 2020; Li et al., 2021; Schuster et al., 2022) is to copy the hidden states of the current token at the exiting layer to all later layers, which will be used to compute the keys and values at later attention layers. Despite its efficiency, this method causes a deviation from the inference process that the model is trained to excel at, which can harm the output quality. Another solution (Corro et al., 2023) is to pre-specify the exiting layer for each token, while ensuring that KV missing in previous tokens will not hinder generation of later tokens; with this approach, the ability of token-wise adaptive selection of exits is inevitably lost. The third method (Bae et al., 2023; Tang et al., 2023) is to store the hidden states of recent tokens that were generated with early exiting, and whenever KV missing happens, run a batch forward pass with the current and recent tokens to fulfill the KV caches. Acceleration by this method relies on the batching effect of GPU computation. In comparison, our proposed pipeline-based method achieves early-exit acceleration in theoretical time complexity.
Other applications of early-exit models.
Interestingly, early exiting has been used for many other purposes in the literature, besides accelerating inference. For example, early-exit losses can provide deep supervision and regularization for training deep networks with enhanced stability of convergence (Szegedy et al., 2015; Lee et al., 2014), which motivates our implementation of the cool-down option for non-constant early-exit loss weights, as explained in Appendix C.1. Other works use early exits for interpreting the intermediate features learned by each layer of a Transformer (Langedijk et al., 2023), for improving outputs of the final exit (Gera et al., 2023), or as the draft models in speculative decoding (Kim et al., 2023b). Early exiting has also found applications in long-tailed classification (Duggal et al., 2020), federated learning with heterogeneous clients (Ilhan et al., 2023; Zhong et al., 2022), token reduction (Liu et al., 2024), layer-wise training of deep networks (Scardapane et al., 2020), among others. We believe that our progress in this work can be beneficial for these various purposes.
Other methods of accelerating LLM inference.
It is worth mentioning that there are other types of dynamic neural networks (Han et al., 2022; Xu & McAuley, 2023) that facilitate conditional computation and elastic inference, such as layer skipping (Bengio et al., 2013, 2015; Wang et al., 2022; Corro et al., 2023; Din et al., 2023; Zeng et al., 2023; Wang et al., 2023) and mixtures of experts (Jacobs et al., 1991; Shazeer et al., 2017; Fedus et al., 2021). Another line of work aims to accelerate LLM inference by designing new decoding methods rather than model architectures (Leviathan et al., 2022; Santilli et al., 2023; Zhao et al., 2023b). Each of these approaches has its own pros and cons, and some of them can be complemented by early exiting. A detailed comparison between these methods is beyond the scope of the current work.